OpenStudy (anonymous):

Simplify i231 where i is the imaginary unit and is defined as: i = sqrt(-1).

OpenStudy (anonymous):

I think the answer would be -1

OpenStudy (anonymous):

yup -1

OpenStudy (anonymous):

no i do not think it is -1

OpenStudy (anonymous):

\[i^{231}\] take the integer remainder when dividing 231 by 4 and you get 3 (since 4 divides 228 evenly) therefore \[i^{231}=i^3=-i\]

OpenStudy (anonymous):

\[i^0=1,i^1=i,i^2=-1,i^3=-i,i^4=1, ...\]

OpenStudy (wasiqss):

answer is indeed -i